# First Self-Renormalized Gluon PDF of Nucleon from Large-Momentum Effective Theory in the Continuum Limit

### v1

@@ Abstract
Lattice QCD provides a first-principles, nonperturbative framework for studying PDFs by discretizing QCD on a spacetime lattice and performing numerical simulations. While traditional global fits rely on experimental cross sections and perturbative QCD factorization, lattice QCD provides a complementary ab initio approach, especially valuable in regions with scarce or uncertain constraints, such as the large-$x$ region ($x\gtrsim 0.4$) relevent for new physics searches and high transverse momentum predictions. Lattice-QCD techniques, such as large-momentum effective theory (LaMET) Ji:2013dva and pseudo-PDF Balitsky:2019krf approaches, enable the extraction of moments and $x$-dependent distributions of many isovector nucleon and valence pion quark PDFs and GPDs; we refer readers interested in learning more about these topics to the reviews in Refs. Ji:2020ect ; Constantinou:2020hdm ; Lin:2023kxn ; Lin:2025hka . In contrast, fewer gluon PDF calculations exist with progress mainly limited to the pseudo-PDF approach: nucleon Fan:2020cpa ; Fan:2022kcb ; HadStruc:2021wmh ; Delmar:2023agv ; HadStruc:2022yaw , pion Good:2023ecp , and kaon Salas-Chavira:2021wui ; NieMiera:2025inn gluon PDFs. Only recently have the first determinations of the nucleon gluon PDFs been achieved using LaMET Good:2024iur ; Good:2025daz .
@@
where we treat $k$, $\Lambda_{\text{QCD}}$, $f(z)$, $d$, $m_{0}$, and $g(z)$ as free fitting parameters. We complete a global fit of these parameters to the interpolated zero momentum $\eta_{s}$ matrix elements for $a\approx\{0.15,0.12,0.09\}$ fm. We use the $\eta_{s}$ data because it has better signal to noise than the nucleon and the operator renormalization should not depend on the external state, and we confirmed that the sum of the linear divergence and renormalon ambiguity are statistically consistent between different hadrons in our previous LaMET study on only one lattice spacing Good:2025daz . We choose $z_{0}=0.05$ fm and $z_{1}=0.36$ fm at a fixed renormalization scale $\mu=2$ GeV, although scale dependence will be explored more carefully in the future.
@@
The global-fit results in $k=0.60(28)$, d=0.003(14)d=-0.003(14), $\Lambda_{\text{QCD}}=0.295(84)$ GeV, and $m_{0}=0.14(31)$ GeV, with a $\chi^{2}/\text{dof}=0.75(48)$, demonstrating that the fit is trustworthy. All of the parameters fall within reasonable ranges as seen in previous quark calculations LatticePartonLPC:2021gpi ; Holligan:2024wpv ; Chen:2024rgi ; Holligan:2023rex . We find that $d$ is statistically consistent with zero likely due to this term being related to subleading logarithms, a fine-tuning parameter. It should also be noted that the parameters are all correlated amongst each other, so the large standard deviations are not as suggestive of poor control over the fits as they may seem. This can be seen in the original self-renormalization paper, which demonstrates that there are valleys of reasonable $\chi^{2}/\text{dof}$ for different values of $k$ and $\Lambda_{\text{QCD}}$, for example LatticePartonLPC:2021gpi .

### v2

@@ Abstract
Lattice QCD provides a first-principles, nonperturbative framework for studying PDFs by discretizing QCD on a spacetime lattice and performing numerical simulations. While traditional global fits rely on experimental cross sections and perturbative QCD factorization, lattice QCD provides a complementary ab initio approach, especially valuable in regions with scarce or uncertain constraints, such as the large-$x$ region ($x\gtrsim 0.4$) relevant for new physics searches and high transverse momentum predictions. Lattice-QCD techniques, such as large-momentum effective theory (LaMET) Ji:2013dva and pseudo-PDF Balitsky:2019krf approaches, enable the extraction of moments and $x$-dependent distributions of many isovector nucleon and valence pion quark PDFs and GPDs; We refer readers interested in learning more about these topics to the reviews in Refs. Ji:2020ect; Constantinou:2020hdm; Lin:2023kxn; Lin:2025hka. In contrast, fewer gluon PDF calculations exist with progress mainly limited to the pseudo-PDF approach: nucleon Fan:2020cpa; Fan:2022kcb; HadStruc:2021wmh; Delmar:2023agv; HadStruc:2022yaw, pion Good:2023ecp, and kaon Salas-Chavira:2021wui; NieMiera:2025inn gluon PDFs. Only recently have the first determinations of the nucleon gluon PDFs been achieved using LaMET Good:2024iur; Good:2025daz.
@@
where we treat $k$, $\Lambda_{\text{QCD}}$, $f(z)$, $d$, $m_{0}$, and $g(z)$ as free fitting parameters. We complete a global fit of these parameters to the interpolated zero momentum $\eta_{s}$ matrix elements for $a\approx\{0.15,0.12,0.09\}$ fm. We use the $\eta_{s}$ data because it has better signal to noise than the nucleon, and the operator renormalization should not depend on the external state, and we confirmed that the sum of the linear divergence and renormalon ambiguity is statistically consistent between different hadrons in our previous LaMET study on only one lattice spacing Good:2025daz. We choose $z_{0}=0.05$ fm and $z_{1}=0.36$ fm at a fixed renormalization scale $\mu=2$ GeV, although scale dependence will be explored more carefully in the future.
@@
The global-fit results in $k=0.61(27)$, d=0.002(10)d=-0.002(10), $\Lambda_{\text{QCD}}=0.286(85)$ GeV, and $m_{0}=0.13(30)$ GeV, with a $\chi^{2}/\text{dof}=0.74(58)$, demonstrating that the fit is trustworthy. All of the parameters fall within reasonable ranges as seen in previous quark calculations LatticePartonLPC:2021gpi; Holligan:2024wpv; Chen:2024rgi; Holligan:2023rex. We find that $d$ is statistically consistent with zero likely due to this term being related to subleading logarithms, a fine-tuning parameter. It should also be noted that the parameters are all correlated with each other, so the large standard deviations are not as suggestive of poor control over the fits as they may seem. This can be seen in the original self-renormalization paper, which demonstrates that there are valleys of reasonable $\chi^{2}/\text{dof}$ for different values of $k$ and $\Lambda_{\text{QCD}}$, for example LatticePartonLPC:2021gpi.